OpenStudy (anonymous):

if there are some parson and dogs are standing in the garden then there are 36 head and 104 legs .find the legs of dogs.please tell with explain, how can we find it

OpenStudy (mathstudent55):

How many legs does one person have? How many heads does one person head?

OpenStudy (mathstudent55):

legs

OpenStudy (mathstudent55):

Also, both a person and a dog have only one leg.

OpenStudy (mathstudent55):

We need to use variables to represent the unknowns. Let's use: p = the number of people d = the number of dogs 1 person has 2 * 1 legs = 2 legs 2 people has 2 * 2 legs = 4 legs 3 people have 2 * 3 legs = 6 legs p people have how many legs?